# A Survey of World Models for Autonomous Driving

### v1

@@ 3.1.4 PC Representations
Early generative efforts targeted realistic LiDAR data simulation for data augmentation and testing. For instance, lidarGeneration [51] adopted a GAN-based framework to mimic real-world LiDAR statistics, while Lidarsim [7] introduced LidarSim to construct near-realistic LiDAR scenes by leveraging empirical sensor distributions. Subsequent works further advanced the diversity and efficiency of point cloud generation. Lidargen [52] proposed a deep learning pipeline that preserves physical consistency while increasing point cloud variability. Meanwhile, AXform [53] employed attention mechanisms to map latent features into 3D space, and PSF [54] developed a straight-flow approach that accelerates the generation process without compromising fidelity. In parallel, NFL [55] introduced Neural LiDAR Fields for synthesizing novel-view point clouds, demonstrating effectiveness under dynamic or partially observed scenes. Extending neural rendering concepts, Nerf-lidar [56] proposed NeRF-LiDAR, leveraging neural radiance fields to produce high-resolution LiDAR data suitable for realistic 3D reconstructions.
A foundational step toward future prediction with PC data is offered by 4DOcc [57], who recast point cloud forecasting as a 4D occupancy prediction problem. Rather than directly inferring raw point coordinates, they map sequential LiDAR scans into intermediate occupancy grids and employ temporal models (e.g., Transformers) to estimate future occupancy states. This paradigm neatly bypasses heavy reliance on meticulously annotated LiDAR data by rendering pseudo LiDAR sequences from the predicted occupancy volumes.
Recent work further demonstrates the synergy of point cloud representations with advanced neural volumetric techniques. For instance, NeMo [46] leverage self-supervised 3D reconstruction and motion cues to achieve robust geometric and dynamic understanding – often outperforming 2D or BEV-based approaches in scenarios with frequent occlusions or varied terrains. Similarly, ViDAR [17] extends this idea by predicting future point cloud distributions from purely visual inputs, thus unifying semantic, spatial, and temporal learning. By integrating generative modeling with latent-space transformations, ViDAR demonstrates notable gains in 3D detection, motion forecasting, and route planning.
@@ 3.2.2 Learning-Based Planning
A central concern in learning-based approaches is balancing performance with safety. SafeDreamer [98] integrates Lagrangian methods into the Dreamer framework to form a safe reinforcement learning algorithm, achieving near-zero safety violations across multiple tasks in the Safety-Gymnasium benchmark. This highlights a growing trend toward incorporating explicit safety constraints into RL-based planners. Another reinforcement learning–based approach, Think2Drive [99], introduces a neural planner capable of handling quasi-realistic scenarios using techniques such as the reset technique, automated scenario generation, termination-priority replay strategy, and steering cost function. This approach demonstrates that an agent can “think” through imagined scenarios in a latent world model to enhance learning efficiency. Tested in CARLA-v2, Think2Drive achieves higher sample efficiency and improved decision quality compared to standard RL baselines.
@@ 4.1 Self-Supervised Learning for World Models
Early work on self-supervised learning for world models in autonomous driving focuses on pure spatio-temporal prediction [104], which estimates future LiDAR scans without explicit tracking. Building on that, UnO [29] and EO [105] introduce occupancy-based representations: UnO uses 4D occupancy fields to model dynamic changes, while EO learns future occupancy by comparing predicted and actual LiDAR sweeps. RenderWorld [106] then integrates multi-view images via a Gaussian-based Img2Occ module to self-supervise 3D occupancy labels, pushing occupancy modeling further by unifying 2D and 3D cues. Subsequently, UniPAD [26] and ViDAR [17] elevate pre-training paradigms: UniPAD harnesses volumetric differentiable rendering for robust 2D–3D representation learning, and ViDAR aligns temporal visual features with point clouds to improve scene understanding and motion forecasting. In parallel, COPILOT4D [107] embraces a discrete diffusion approach to generate future states from tokenized sensor data, capturing complex distributions without labeled samples. For end-to-end driving, SSR [108] introduces a sparse scene representation that leverages only a handful of navigation tokens with a self-supervised temporal alignment module, while CarFormer [35] utilizes slot attention in a BEV framework to learn object-centric representations. This end-to-end trend is further advanced by the LAW [109], which predicts future latent features conditioned on ego actions to reduce supervision needs. Finally, BEVWorld [27] unifies diverse sensor modalities within a single BEV latent space using diffusion modeling, showcasing the current push toward holistic, self-supervised world models that fuse complementary signals for perception, prediction, and planning in autonomous driving.
@@ 5.1 Scene Understanding
Scene understanding serves as the foundation for real-time environmental perception and comprehension in autonomous driving. This technology enables autonomous systems to identify roads, vehicles, pedestrians, and traffic signs, thereby facilitating safe and effective driving decisions. With the advancement of the field, world models are introduced to enhance the depth and breadth of scene understanding. By integrating data from various sensors, world models build 3D representations of the environment and predict future scene evolution, bridging information gaps while improving safety and efficiency.
@@ 5.3 Simulation
Simulation plays a pivotal role in autonomous driving by providing controlled environments that can test diverse scenarios without real-world risks. However, traditional simulation methods often fail to capture the full complexity of dynamic driving conditions, prompting researchers to explore world models – comprehensive frameworks that encompass both spatial and temporal aspects of the driving context. For instance, the OG Representation-based OccSora [34] is a diffusion-based 4D occupancy generation model designed to simulate the evolution of 3D scenes for autonomous driving. OccSora employs a 4D scene tokenizer to extract compact spatio-temporal representations, thereby enabling high-fidelity reconstruction of extended occupancy videos. By learning a diffusion transformer on these representations, OccSora generates 16-second videos with realistic 3D layouts and temporal consistency, demonstrating an understanding of the spatial and temporal distributions in driving scenes. This trajectory-aware 4D generation serves as a world simulator to inform decision-making in autonomous vehicles. Another method, image Representation-based SimGen [113] learns to produce diverse driving scenes by integrating data from both simulators and the real world. Through a cascade diffusion pipeline, SimGen bridges the gap between simulated and real-world data by adhering to simulator-driven layout guidance and rich text prompts, ultimately yielding realistic driving scenarios. By combining simulated and real-world data, SimGen boosts the diversity and authenticity of generated scenes, which is essential for training robust autonomous driving systems. These advancements in world model-driven simulation highlight the potential for creating more realistic and adaptable virtual environments, thereby strengthening the robustness and reliability of autonomous driving solutions.
@@ 6.4 Planning in Driving Scenarios: Motion planning
Results. Table III presents a quantitative comparison of motion planning methods on the nuScenes [122] dataset, encompassing various sensor inputs (LiDAR, camera, and 3D occupancy) alongside different levels of auxiliary supervision (maps, bounding boxes, etc.). End-to-end autonomous driving frameworks (e.g., UniAD [134]) display strong results in both trajectory accuracy and collision avoidance, especially when trained on rich annotations (map, box, and motion supervision). By contrast, occupancy-driven methods (e.g., OccWorld [47], OccNet [136], RenderWorld [106], DFIT-OccWorld [50]) reduce reliance on auxiliary data, yet remain competitive in purely camera-based scenarios – indicating robust performance under more constrained conditions.
@@ 7.1 Self-Supervised World Models
Reducing Label Dependency. Future efforts may explore coupling SSL with decoupled dynamic-flow mechanisms or voxel deformation approaches (e.g., DFIT-OccWorld [50]) to learn high-level scene dynamics without explicit ground-truth annotations, echoing the occupancy-centric strategies in UnO [29] and EO [105]. In addition, large-scale generative tasks such as video prediction (InfinityDrive [28]) or multimodal reconstruction (CarFormer [35]) can serve as powerful self-supervised objectives that capture complex 4D structures and reduce susceptibility to labeling errors – mirroring the rendering-based paradigm in RenderWorld [106] and the integrated BEV approach in BEVWorld [27].
Exploring Unlabeled Data Potential. Emerging methods like Think2Drive [99] and Symphony [137] illustrate how reinforcement learning (RL) agents and generative models can leverage vast unlabeled or partially labeled datasets to discover underlying spatiotemporal structures. By fusing simulation-based “thinking” with real-world sensor data, future frameworks could build on the latent forecasting strategies in LAW [109] or the temporal alignment techniques in ViDAR [17], thereby capturing subtle dynamics – from minor lane deviations to complex urban interactions – in a self-supervised manner. Such advances would not only reduce the cost of annotation but also sharpen domain-relevant representations, paving the way for safer and more robust autonomous driving systems.

### v2

@@ 3.1.4 PC-based Generation
PC-based Generation predicts future point clouds from past point cloud inputs, and existing methods [71, 72] directly forecast future laser points despite challenges posed by sparse and irregular LiDAR data. PointRNN [71] employs a recurrent neural network to capture spatiotemporal features in moving point clouds by incorporating point-level operations, while MoNet [72] leverages motion information to model dynamic variations in continuous point cloud sequences through embedding motion features into its network architecture. Another line of methods uses range images [73, 74], a representation obtained by projecting point clouds to dense 2D images using sensor intrinsic and extrinsic parameters. Based on historical range images, they apply 3D convolutions [75, 76] or LSTMs [77, 78] to predict future point clouds, yet they additionally model the motion of sensor intrinsic and extrinsic parameters.
On the one hand, PC-based Generation have gradually evolved to target realistic LiDAR data simulation for data augmentation and testing. SE3-Nets [79] predicts a scene point cloud in the next frame by considering the action applied to an object in the scene, while lidarGeneration [80] uses VAEs and GANs to generate a scene point cloud from a random noise vector, and Dscnet [81] generates a high-resolution scene point cloud conditioned on stereo images and low-resolution LiDAR inputs. lidarGeneration [80] adopted a GAN-based framework to mimic real-world LiDAR statistics, whereas Lidarsim [7] constructs near-realistic LiDAR scenes by leveraging empirical sensor distributions. Subsequent works have further advanced the diversity and efficiency of point cloud generation: Lidargen [82] proposed a deep learning pipeline that preserves physical consistency while increasing point cloud variability, AXform [83] employed attention mechanisms to map latent features into 3D space, and PSF [84] developed a straight-flow approach that accelerates the generation process without compromising fidelity. In parallel, NFL [85] introduced Neural LiDAR Fields for synthesizing novel-view point clouds, demonstrating effectiveness under dynamic or partially observed scenes, and extending neural rendering concepts, Nerf-lidar [86] proposed NeRF-LiDAR, which leverages neural radiance fields to produce high-resolution LiDAR data suitable for realistic 3D reconstructions.
On the other hand, Recent advancements in point cloud generation have increasingly intertwined with world models for autonomous driving. 4DOcc [87] transforms sequential LiDAR scans into intermediate occupancy grids and utilizes temporal models, such as Transformers, to predict future occupancy states, thereby reducing dependence on extensively annotated LiDAR data. Copilot4D [88] tokenizes LiDAR observations using a Vector Quantized Variational Autoencoder (VQVAE) and forecasts future point clouds through discrete diffusion, enabling unsupervised learning of world models. NeMo [89] employs self-supervised 3D reconstruction and motion cues to achieve a robust understanding of geometry and dynamics, often surpassing 2D or Bird’s-Eye View (BEV)-based methods in complex environments. ViDAR [18] introduces a pre-training framework for autonomous driving by predicting future 3D point cloud sequences from historical visual inputs, effectively integrating semantic, spatial, and temporal learning.
@@ 3.2.2 Learning-Based Planning
Model-based reinforcement learning (MBRL) leverages world models to predict environmental states, improving data efficiency and making it suitable for autonomous driving due to high sample efficiency and learnable state transitions [128]. Early attempts like MILE [32] introduced model-based approaches to autonomous driving, though their reliance on expert-collected data limited performance [128]. MBOP [129] comprehensively addresses planning challenges by utilizing multiple offline-learned models within a model-predictive control (MPC) framework, enabling flexible reward function extensions and state constraint incorporation. While enhancing interpretability through learned dynamics models, MBOP employs simplistic deterministic models that ignore environmental stochasticity and aleatoric uncertainty [128]. UMBRELLA [130] advances this paradigm by employing stochastic dynamics models to capture diverse traffic scene evolutions, explicitly addressing both epistemic and aleatoric uncertainties while learning from offline data. This approach integrates partial observability considerations and uses interpretable representations to tackle simultaneous prediction, planning, and control challenges in self-driving vehicles (SDVs). SafeDreamer [131] enhances safety considerations by integrating Lagrangian methods into the Dreamer framework, achieving near-zero safety violations in Safety-Gymnasium benchmarks. This reflects the growing emphasis on embedding explicit safety constraints into reinforcement learning (RL)-based planners for real-world applications. Think2Drive [128] models environmental transitions through world models [12], employing them as neural network simulators for latent-space planning. The method introduces a neural planner with reset techniques, automated scenario generation, and steering cost functions, demonstrating how agents can improve learning efficiency by ”thinking” through imagined scenarios in latent world models.
@@ 4.1 Self-Supervised Learning for World Models
Early work on self-supervised learning for world models in autonomous driving focuses on pure spatio-temporal prediction [75], which estimates future LiDAR scans without explicit tracking. Building on that, UnO [30] and EO [60] introduce occupancy-based representations: UnO uses 4D occupancy fields to model dynamic changes, while EO learns future occupancy by comparing predicted and actual LiDAR sweeps. RenderWorld [70] then integrates multi-view images via a Gaussian-based Img2Occ module to self-supervise 3D occupancy labels, pushing occupancy modeling further by unifying 2D and 3D cues. Subsequently, UniPAD [27] and ViDAR [18] elevate pre-training paradigms: UniPAD harnesses volumetric differentiable rendering for robust 2D–3D representation learning, and ViDAR aligns temporal visual features with point clouds to improve scene understanding and motion forecasting. In parallel, COPILOT4D [135] embraces a discrete diffusion approach to generate future states from tokenized sensor data, capturing complex distributions without labeled samples. For end-to-end driving, SSR [136] introduces a sparse scene representation that leverages only a handful of navigation tokens with a self-supervised temporal alignment module, while CarFormer [52] utilizes slot attention in a BEV framework to learn object-centric representations. This end-to-end trend is further advanced by the LAW [51], which predicts future latent features conditioned on ego actions to reduce supervision needs. Finally, BEVWorld [28] unifies diverse sensor modalities within a single BEV latent space using diffusion modeling, showcasing the current push toward holistic, self-supervised world models that fuse complementary signals for perception, prediction, and planning in autonomous driving. AD-L-JEPA [137] introduces a self-supervised framework for autonomous driving using LiDAR data, leveraging a Joint Embedding Predictive Architecture (JEPA) to learn spatial world models by predicting BEV embeddings, eliminating generative/contrastive mechanisms and explicit data reconstruction while capturing occluded or uncertain environmental details. AD-L-JEPA achieves 5×5$\times$5 × faster pre-training than SOTA approaches (e.g., Occupancy-MAE [138], ALSO [139]) by avoiding contrastive pair curation and generative overhead, while demonstrating robust transfer learning even with partially randomized encoder initialization and superior label efficiency on downstream tasks like 3D object detection.
@@ 5.1 Scene Understanding
Scene understanding serves as the foundation for real-time environmental perception and comprehension in autonomous driving. This technology enables autonomous systems to identify roads, vehicles, pedestrians, and traffic signs—thereby facilitating safe and effective driving decisions. With the advancement of the field, world models are introduced to enhance the depth and breadth of scene understanding. By integrating data from various sensors, world models build 3D representations of the environment and predict future scene evolution, bridging information gaps while improving safety and efficiency.
@@ 5.3 Simulation
Simulation plays a pivotal role in autonomous driving by providing controlled environments that can test diverse scenarios without real-world risks. However, traditional simulation methods often fail to capture the full complexity of dynamic driving conditions, prompting researchers to explore world models – comprehensive frameworks that encompass both spatial and temporal aspects of the driving context. For instance, modern platforms like CARLA [148] (Unreal Engine-based) and AirSim [149] (Microsoft’s aerial/ground simulator) leverage game engines to synthesize high-fidelity environments with customizable sensors (LiDAR, cameras), dynamic weather, and traffic scenarios. While these tools enable physics-based interactions and modular testing, their reliance on synthetic data introduces a reality gap: overly perfect textures/lighting and scripted traffic patterns fail to capture natural imperfections or rare edge cases. Recently, another line of simulations, which are based on world models, is gaining ground. The OG Representation-based OccSora [35], is a diffusion-based 4D occupancy generation model designed to simulate the evolution of 3D scenes for autonomous driving. OccSora employs a 4D scene tokenizer to extract compact spatio-temporal representations, thereby enabling high-fidelity reconstruction of extended occupancy videos. By learning a diffusion transformer on these representations, OccSora generates 16-second videos with realistic 3D layouts and temporal consistency, demonstrating an understanding of the spatial and temporal distributions in driving scenes. This trajectory-aware 4D generation serves as a world simulator to inform decision-making in autonomous vehicles. In contrast, image Representation-based SimGen [50] learns to produce diverse driving scenes by integrating data from both simulators and the real world. Through a cascade diffusion pipeline, SimGen bridges the gap between simulated and real-world data by adhering to simulator-driven layout guidance and rich text prompts, ultimately yielding realistic driving scenarios. By combining simulated and real-world data, SimGen boosts the diversity and authenticity of generated scenes, which is essential for training robust autonomous driving systems. These advancements in world model-driven simulation highlight the potential for creating more realistic and adaptable virtual environments, thereby strengthening the robustness and reliability of autonomous driving solutions.
@@ 6.4 Planning in Driving Scenarios: Motion planning
Results. Table III presents a quantitative comparison of motion planning methods on the nuScenes [151] dataset, encompassing various sensor inputs (LiDAR, camera, and 3D occupancy) alongside different levels of auxiliary supervision (maps, bounding boxes, etc.). End-to-end autonomous driving frameworks (e.g., UniAD [162]) display strong results in both trajectory accuracy and collision avoidance, especially when trained on rich annotations (map, box, and motion supervision). By contrast, occupancy-driven methods (e.g., OccWorld [64], OccNet [164], RenderWorld [70], DFIT-OccWorld [65]) reduce reliance on auxiliary data, yet remain competitive in purely camera-based scenarios—indicating robust performance under more constrained conditions.
@@ 7.1 Self-Supervised World Models
Reducing Label Dependency. Future efforts may explore coupling SSL with decoupled dynamic-flow mechanisms or voxel deformation approaches (e.g., DFIT-OccWorld [65]) to learn high-level scene dynamics without explicit ground-truth annotations, echoing the occupancy-centric strategies in UnO [30] and EO [60]. In addition, large-scale generative tasks such as video prediction (InfinityDrive [29]) or multimodal reconstruction (CarFormer [52]) can serve as powerful self-supervised objectives that capture complex 4D structures and reduce susceptibility to labeling errors—mirroring the rendering-based paradigm in RenderWorld [70] and the integrated BEV approach in BEVWorld [28].
Exploring Unlabeled Data Potential. Emerging methods like Think2Drive [128] and Symphony [165] illustrate how reinforcement learning (RL) agents and generative models can leverage vast unlabeled or partially labeled datasets to discover underlying spatiotemporal structures. By fusing simulation-based “thinking” with real-world sensor data, future frameworks could build on the latent forecasting strategies in LAW [51] or the temporal alignment techniques in ViDAR [18], thereby capturing subtle dynamics—from minor lane deviations to complex urban interactions—in a self-supervised manner. Such advances would not only reduce the cost of annotation but also sharpen domain-relevant representations, paving the way for safer and more robust autonomous driving systems.